INAF queue observing, 2020 Jan 28 - Jan 29

Observers: A. Rossi, S. Paiano
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: MODS1-2

At sunset the sky is clear Perfect night!

Summary

Night duration (12 degree - twilight) 1:30 - 13:30 UT = 12h 00m = 43200 seconds

Observing time: 11h 15min (93%)

Weather time loss 20min (2%) 1200sec

Technical time loss: 35min (5%) 2100sec

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibrations   MODS1-2 5as dual grating     for STD dual
Calibrations   MODS1-2 arcs redgrating     for Belladitta
Calibrations   MODS1-2 slitflat 1.2" redgrating     for Belladitta
Calibrations   MODS1-2 ugri flat ima     for Rossi, z was not completed
STD GD71 MODS1-2 dual and red grating   mods1r.0031-0034,mods1b.0019-20 mods2r.0033-36,mods2b.0019-20 for Rossi, z was not completed
Mannucci J0415 MODS1-2 dual grating, 0.6" 0.5/0.84 mods1r.0028,39,54 mods1b.0018,21,26 mods2r.0029,41,54 mods2b.0018,21,26 3 pointings
Rossi Sn2018ibb MODS1-2 dual grating, 1.2" 1/3h mods1r.0043-0044,mods1b.0022-23 mods2r.0046-47,mods2b.0022-23 completed
Rossi Sn2018ibb MODS1-2 ugri+ugrz imaging 0.2h mods1r.0045-0046,mods1b.0024-25 mods2r.0024-25,mods2b.0048-50 completed
Wolter NGC2276 S6/3-c MODS1-2 dual grating 2h mods1r.0056-61 ,mods1b.0027-32 mods2r.0060-65,mods2b.0027-32 for A.Wolter CHECK
Wolter NGC2276 S6/3-c MODS1-2 g+r imaging 240 sec mods1r.0062- ,mods1b.0033- mods2r.0066-,mods2b.0033- for A.Wolter CHECK
Paiano 4FGLJ0955+35 MODS1-2 dual grating 1.2" 3h mods1r.0083-89 ,mods1b.0043-49- mods2r.0087-93,mods2b.0044-50  
Paiano GD153 MODS1-2 dual grating 5"   mods1r.0092-94 ,mods1b.0050-52- mods2r.0097-99,mods2b.0051-53 SDT for Paiano DDT
Paiano GD153 MODS1-2 dual grating 1.2"   mods1r.0095-97 ,mods1b.0053-55- mods2r.0100-102,mods2b.0054-56 SDT for Paiano DDT
Paiano 4FGLJ0955+35 MODS1-2 dual grating 1.2" 0.6h mods1r.0104-105 ,mods1b.0056-0057- mods2r.0108-0109-,mods2b.0057-0058 completed

Details

01:34 Slew to Mannucci program, target J0415
  • PA -30 ok for -2<HA<0
  • DIMM seeing 0.8"
  • acq mods1r.20200129.0025-0027, mods2r.20200129.0026-0028
  • sci mods1r.0028,mods1b.0018 mods2r.0029,mods2b.0018- just 1 300sec

*01:57 Slew to STD GD71 dual and red grating for Rossi, Belladitta, Paiano, Wolter
  • acq mods1r.20200129.0029-0030, mods2r.20200129.0030-0032
  • sci mods1r.0031-0034,mods1b.0019-20 mods2r.0033-36,mods2b.0019-20 dual and red grating

02:27 Slew to Mannucci program, target J0415
  • PA -30 ok for -2<HA<0
  • DIMM seeing 1.0"
  • acq mods1r.0035-0038, mods2r.0037-0040 1 more because offset adjusted manually
  • sci mods1r.0039,mods1b.0021 mods2r.0041,mods2b.0021 just 1 300sec

02:51 Slew to Rossi DDT target SN2018ibb, PA +10
  • DIMM seeing 0.8"
  • acq mods1r.0040-0042, mods2r.0042-0045
  • sci mods1r.0043-0044,mods1b.0022-23 mods2r.0046-47,mods2b.0022-23
  • sci mods1r.0045-0046,mods1b.0024-25 mods2r.0024-25,mods2b.0048-50 imaging ug+ri and ug+rz

03:57 Slew to Mannucci program, target J0415
  • PA 220 ok for 0<HA<2
  • DIMM seeing 0.8"
  • acq mods1r.0047-0051, mods2r.0050-0053
  • sci mods1r.0052,mods1b.0026 mods2r.0054,mods2b.0026 just 1 300sec

04:22 Slew to Wolter program NGC2276 S6/3-c
  • DIMM seeing 0.8"
  • acq mods1r.0053-0055, mods2r.0055-0059
  • sci mods1r.0056-61 ,mods1b.0027-32 mods2r.0060-65,mods2b.0027-32

06:05 We continue the Wolter's program NGC2276 S6/3-c with imaging
  • DIMM seeing 0.73"
  • The 50'' is too big, lost guider, and hit the copointing limit.
  • We create a new OB with acquisition on 50'' offset and take the imaging
  • acq =mods1r.00 =, =mods2r.00 =
  • sci mods1r.0062-63 ,mods1b.0033-34- mods2r.0066-67-,mods2b.0033-34-

06:37 We are continuing the Wolter's program NGC2276 S6/3-c providing the spectra with offset +-5" in x (left and right).
  • DIMM seeing 1.00"
  • acq mods1r.0064-0068, mods2r.0068-72
  • sci-ima mods1r.0064-0068, mods2r.0068-72
  • = 10 min loss= During the acquisition, we had a problem with the IMCS-Lock of MODS2 and we lost an acquisition slit image. We re-did the procedure and the second time every acquisition images was taken.
  • sci-spec-Cielo1 mods1r.0071-73 ,mods1b.0037-39- mods2r.0075-77,mods2b.0037-40
  • We did also an imaging without offset before to take the spectra.
  • 10 min loss mods2 red frozen at second exposure
  • DIMM seeing 0.6"
  • sci-spec-Cielo2 mods1r.0074- ,mods1b.0040- mods2r.0078-,mods2b.0041-

08:36 We are observing the target 4FGLJ0955+35 of the Paiano DDT.
  • DIMM seeing 0.95"
  • acq =mods1r.0077-0082 =, =mods2r.0081-86 =
  • DIMM seeing 0.8"
  • Problem with the guide stars: the stars are different for the two telescopes. This happened because there is another star with similar magnitude close the guide star choosen. We lost about 15 minutes to fix it.
  • sci mods1r.0083-89 ,mods1b.0043-49- mods2r.0087-93,mods2b.0044-50
  • Weather is perfect: T=-3C. Humiditu=32%, Wind=8m/s, no clouds!
  • DIMM seeing 0.9" and stable
  • The MODS-1 B spectrum number 0047 has to be rejected because is noise - we re-did the observation to obtain this missing frame.

10:40 We observe the standard star GD153 for the Paiano DDT with slit 5
  • DIMM seeing 0.98"
  • acq =mods1r.0090-91 =, =mods2r.0094-96 =
  • DIMM seeing 1.0"
  • sci mods1r.0092-94 ,mods1b.0050-52- mods2r.0097-99,mods2b.0051-53

11:00 We observe the standard star GD153 for the Paiano DDT with slit 1.2
  • DIMM seeing 0.98"
  • acq skipped because we use the previous one, although it is not optimal
  • DIMM seeing 1.0"
  • sci mods1r.0095-97 ,mods1b.0053-55- mods2r.0100-102,mods2b.0054-56

11:20 At the moment to change the source, some clouds were approaching. We stopped the observations for 10 minutes. No sources were available for the observations due to the faintness of the sources, the sky position contrains and seeing restriction. We decided to re-repoint again 4FGL0955+35 that can be observed also with no-spectrophotometric sky.
  • DIMM seeing 1.00"
  • acq =mods1r.0098-102 =, =mods2r.00103-107 =
  • sci mods1r.0104-105 ,mods1b.0056-0057- mods2r.0108-0109-,mods2b.0057-0058

Notes from the night: We has IMCS lock problems all night and in total we estimate around 10 min of lost time

13:40 Ciao, we close the shop!
Topic revision: r10 - 07 Feb 2020, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback